OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point shown and is perpendicular to the line shown.

OpenStudy (goformit100):

y=mx+c

OpenStudy (anonymous):

this is the graph

OpenStudy (anonymous):

Y-2=-1/3(x-6) 3y-6=-x+6 3y=-x+12

OpenStudy (anonymous):

thank u